# SENSEI: Semantic Exploration Guided by Foundation Models to Learn Versatile World Models

### v1

@@ Abstract
Exploration is a cornerstone of reinforcement learning (RL). Intrinsic motivation attempts to decouple exploration from external, task-based rewards. However, established approaches to intrinsic motivation that follow general principles such as information gain, often only uncover low-level interactions. In contrast, children’s play suggests that they engage in meaningful high-level behavior by imitating or interacting with their caregivers. Recent work has focused on using foundation models to inject these semantic biases into exploration. However, these methods often rely on unrealistic assumptions, such as language-embedded environments or access to high-level actions. We propose SEmaNtically Sensible ExploratIon (Sensei), a framework to equip model-based RL agents with an intrinsic motivation for semantically meaningful behavior. Sensei distills a reward signal of interestingness from Vision Language Model (VLM) annotations, enabling an agent to predict these rewards through a world model. Using model-based RL, Sensei trains an exploration policy that jointly maximizes semantic rewards and uncertainty. We show that in both robotic and video game-like simulations Sensei discovers a variety of meaningful behaviors from image observations and low-level actions. Sensei provides a general tool for learning from foundation model feedback, a crucial research direction, as VLMs become more powerful. [^1] Project website: https://sites.google.com/view/sensei-paper
@@ 1 Introduction
A potential solution in the age of Large Language Models (LLMs), is to utilize language as a cultural-transmitter to inject “human notions of interestingness" (Zhang et al., 2023a) into RL agents’ exploration. LLMs are trained on an immense amount of data produced mostly by humans. Thus, their responses are likely to mirror preferences about what humans find interesting. However, the most prominent works in this domain assume (1) a language-grounded environment (Zhang et al., 2023b; Du et al., 2023; Klissarov et al., 2023), (2) the availability of an offline dataset with exhaustive state-space coverage (Klissarov et al., 2023), or (3) access to high-level actions (Zhang et al., 2023a; Du et al., 2023). These assumptions are still detached from the current reality of embodied agents, e.g. in robotics, which don’t come with perfect state or event captioners, pre-existing offline datasets nor with robust, high-level actions. Furthermore, none of these approaches learn an internal model of “interestingness.” Thus, they rely on the LLM, or a distilled module, to continuously guide their exploration and fail to transfer this knowledge to novel states.
In this work, we propose SEmaNtically Sensible ExploratIon (Sensei), a framework for Vision Language Model (VLM) guided exploration for model-based RL agents, illustrated in Fig. 1. Sensei starts with a short description of the environment and a dataset of observations (e.g. images) collected through self-supervised exploration. A VLM is prompted to compare the observations pairwise with respect to their interestingness and the resulting ranking is distilled into a reward function. When the agent explores its environment, it receives semantically-grounded exploration rewards from Sensei. It learns to predict this exploration signal through its learned world model, corresponding to an internal model of “interestingness”. The agent improves its exploration strategy by aiming for states for which it predicts a high interestingness and then branching out to uncertain situations.
Our main contributions are as follows:
@@ 2 Method
Figure 2: Intrinsic rewards in Sensei: (a) Prior to task-free exploration, we prompt GPT-4 to compare images with respect to the interestingness for a certain environment. From the resulting ranking we distill a reward function $R_{\psi}$ using VLM-Motif. (b) Later, an agent learns an Rssm world model from task-free exploration. From each model state, the agent predicts different sources of intrinsic rewards, i.e., epistemic uncertainty-based reward and our distilled semantic reward.
2.1 Reward function distillation: MOTIFate your SENSEI
@@ 2.3 Exploration policy: Go and Explore with SENSEI
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: Semantic exploration rewards for example trajectories: From top to bottom we show example trajectories for Minihack KeyRoom-S15 and KeyChest (see Fig. 9 for map views) and two Robodesk episodes. We showcase rewards from VLM-Motif distilled from GPT-4 annotations using Plan2Explore (P2X) data. The reward trajectories peak at the “interesting” moments of exploration, such as opening a drawer in Robodesk or picking up the key in MiniHack. For Robodesk we show reward trajectories for both Sensei and a version of Sensei with a more general, zero pre-knowledge prompting strategy (Sensei general, see Sec. 4.2.2)
@@ 3 Related work
Intrinsic rewards are applied either to facilitate exploration in tasks where direct rewards are sparse or in a task-agnostic setting where they help collect diverse data. There are many different reward signals that have been proposed as exploration rewards (Baldassarre & Mirolli, 2013), such as prediction error (Schmidhuber, 1991; Pathak et al., 2017; Kim et al., 2020), Bayesian surprise (Storck et al., 1995; Blaes et al., 2019; Paolo et al., 2021), learning progress (Schmidhuber, 1991; Colas et al., 2019; Blaes et al., 2019), empowerment (Klyubin et al., 2005; Mohamed & Jimenez Rezende, 2015), metrics for state-space coverage (Bellemare et al., 2016; Tang et al., 2017; Burda et al., 2019) and regularity (Sancaktar et al., 2024). While effective for low-dimensional observations, such objectives are challenging to apply for high-dimensional image observations. Here, alternatives are employing low-dimensional goal spaces (Colas et al., 2019; OpenAI et al., 2021; Nair et al., 2018; Pong et al., 2019; Zadaianchuk et al., 2021; Mendonca et al., 2021) or learning latent world models (Hafner et al., 2019a, 2023; Gumbsch et al., 2024) that can be employed for model-based exploration (Pathak et al., 2019; Sekar et al., 2020). In particular, Plan2Explore (Sekar et al., 2020) uses ensemble disagreement of latent space dynamics predictions as an intrinsic reward. While this is a very general strategy for exploration, this could be limited in more challenging environments where semantically meaningful or goal-directed behavior (Spelke, 1990) is needed for efficient exploration.
@@
Exploration with foundation models: Recent improvements of in-context learning of LLMs open additional ways to explore using human bias of interestingness during exploration (Klissarov et al., 2023; Du et al., 2023; Zhang et al., 2023a) and skill learning (Colas et al., 2020, 2023; Zhang et al., 2023b). Motif (Klissarov et al., 2023) leverages LLMs to derive intrinsic rewards by comparing pairs of event captions, demonstrating its efficacy in the complex game of NetHack (Küttler et al., 2020). Similarly, ELLM (Du et al., 2023) uses LLMs to guide RL agents towards goals that are meaningful, based on the agent’s current state represented by text, showing improved task coverage in the Crafter environment (Hafner, 2021). Furthermore, OMNI (Zhang et al., 2023a) introduces a novel method to prioritize tasks by modeling human notions of interestingness using LLMs. Thereby, OMNI focuses on tasks that are not only learnable but also generally interesting. LAMP (Adeniji et al., 2023) proposes to use VLMs for reward modulation. First, a set of potential tasks are generated with an LLM and then LAMP uses VLMs to generate rewards for these tasks to learn a language-conditioned policy. This policy is later finetuned with actual task rewards.
@@
Reward-shaping through VLMs: Most works that rely on VLMs as reward sources try to solve the reward specification problem in RL. In these works, a task is assumed to be described as a language caption (Cui et al., 2022; Rocamonde et al., 2023; Baumli et al., 2023; Adeniji et al., 2023), as a goal image (Cui et al., 2022), or as a video demonstration (Sontakke et al., 2023). In particular, RL-VLM-F (Wang et al., 2024) uses a very similar setup to Sensei. Pairs of images from initial rollouts are compared using a VLM to distill a reward function via Motif (Klissarov et al., 2023). However, unlike in our work, the VLM is explicitly prompted with the task, whereas we distill an environment-specific but general exploration reward. Furthermore, Sensei assumes a model-based setup to learn a world model instead of optimizing a task-based policy.
@@ 4 Results
We answer these questions by (1) illustrating that the semantic rewards obtained from VLM-Motif reflect interesting events in the environment, (2) quantitatively showing that Sensei leads to more interaction-rich behavior during task-free exploration, and (3) employing the learned world models to successfully train task-based policies. We use two fundamentally different types of environments:
@@
We collect the initial dataset $\mathcal{D}^{\mathrm{init}}$ with Plan2Explore (Sekar et al., 2020), the current state-of-the art in exploration with pixel-based observations. We collect data from 500k steps of exploration in Minihack and 1M steps in Robodesk. For data annotation, we use GPT-4 (details in Suppl. C.3).
@@ 4.1 Reward function of SENSEI
We illustrate how the distilled VLM-Motif reward function $R_{\psi}$ assigns semantic exploration rewards $r^{\mathrm{sem}}_{t}$ for exemplary sequences from MiniHack and Robodesk, shown in Fig. 4. In MiniHack, we clearly see jumps in $r^{\mathrm{sem}}_{t}$ for significant events. Frames 2 & 3 in KeyRoom-S15 and KeyChest respectively, are right before the key is picked up. Later, $r^{\mathrm{sem}}_{t}$ increases further once the agent is at the door or chest with a key (Frame 3 in KeyRoom-S15 and Frames 4&5 in KeyChest). For Robodesk, we see that as the robot is interacting with objects, $r^{\mathrm{sem}}_{t}$ increases for the examples of opening the drawer and pushing the blocks.
@@ 4.2.2 Robodesk
Next, we analyze exploration in the challenging visual control suite of Robodesk. Here we compare 1M steps of exploration in Sensei with Plan2Explore and Random Network Distillation (RND, Burda et al., 2019), a strong model-free exploration approach that uses prediction errors of random image embeddings as intrinsic rewards to maximize state space coverage. Fig. 6 plots the mean number of object interactions during exploration for the three methods. On average, Sensei interacts more with most available objects than the baselines. As a result, in a majority of tasks Sensei receives more task rewards during exploration than Plan2Explore or RND (shown in Suppl. D.3). Qualitatively, we observe that Plan2Explore mostly performs arm stretches [^3] Interestingly, this can still lead to solving tasks. For example, stretching the arm against the sliding cabinet can close it, and stretching the arm toward a block can push it off the table., whereas RND mostly moves the arm around in the center of the screen, mostly hitting buttons, as they are also centered on the table, and occasionally hitting objects.
Thus, our semantic exploration reward seems to lead to more meaningful behavior than pure epistemic uncertainty-based exploration, even in a low-level motor control robotic environment.
@@
In the previous Sensei experiments, we provided a small environment description in the prompts for the VLM annotations. We investigate whether Sensei relies on this external description in Robodesk, and compare against a version of Sensei using a more general, zero-knowledge prompting strategy (Sensei general). Sensei general first prompts the VLM for an environment description given an image of the environment and uses the generated answer as context to annotate the dataset of preferences (details in Suppl. C.3.2). As shown in Fig. 6, Sensei general interacts roughly as often with the relevant objects as Sensei, outperforming both Plan2Explore and RND in terms of overall number of object interactions. Thus, injecting external environment knowledge to the prompts is not necessary and this step can be fully automated. This further cements the generality of our approach.
@@
We perform ablations to see (1) how noisy annotations from VLMs affect Sensei compared to an oracle annotator, (2) how much the behavior richness of the initial dataset affects Sensei’s performance (Suppl. Fig. 12), and (3) ablate our Go-explore switching strategy. We observe that as VLMs get better, there is indeed more to gain from Sensei, and richer exploration data helps Sensei bootstrap faster. See Suppl. D.2 for more information. We further showcase the robustness of our Go-Explore switching strategy in terms of hyperparameter sensitivity compared to a variant of Sensei where the semantic and disagreement rewards get fixed weights (Suppl. D.6).
@@ 4.3 Fast downstream task learning
(a) extrinsic phase in KeyRoom-S15
(b) extrinsic phase in KeyChest
Figure 7: Downstream task performance in MiniHack: We plot the mean of the episode score obtained during evaluation for the MiniHack tasks (a) KeyRoom-S15 and (b) KeyChest, with world models learned from Sensei vs. Plan2Explore (P2X) exploration. We also show results for learning a task policy from scratch with DreamerV3, and the model-free baseline PPO. Shaded areas depict the standard error (10 seeds) and we apply smoothing over the score trajectories with window size 3.

### v2

@@ Abstract
Exploration is a cornerstone of reinforcement learning (RL). Intrinsic motivation attempts to decouple exploration from external, task-based rewards. However, established approaches to intrinsic motivation that follow general principles such as information gain, often only uncover low-level interactions. In contrast, children’s play suggests that they engage in meaningful high-level behavior by imitating or interacting with their caregivers. Recent work has focused on using foundation models to inject these semantic biases into exploration. However, these methods often rely on unrealistic assumptions, such as language-embedded environments or access to high-level actions. We propose SEmaNtically Sensible ExploratIon (Sensei), a framework to equip model-based RL agents with an intrinsic motivation for semantically meaningful behavior. Sensei distills a reward signal of interestingness from Vision Language Model (VLM) annotations, enabling an agent to predict these rewards through a world model. Using model-based RL, Sensei trains an exploration policy that jointly maximizes semantic rewards and uncertainty. We show that in both robotic and video game-like simulations Sensei discovers a variety of meaningful behaviors from image observations and low-level actions. Sensei provides a general tool for learning from foundation model feedback, a crucial research direction, as VLMs become more powerful. [^1] Project website with videos and code: https://sites.google.com/view/sensei-paper
@@ 1 Introduction
A potential solution in the age of Large Language Models (LLMs), is to utilize language as a cultural-transmitter to inject “human notions of interestingness" (Zhang et al., 2023a) into RL agents’ exploration. LLMs are trained on an immense amount of data produced mostly by humans. Thus, their responses are likely to mirror human preferences. However, the most prominent works in this domain assume (1) a language-grounded environment (Zhang et al., 2023b; Du et al., 2023; Klissarov et al., 2023), (2) the availability of an offline dataset with exhaustive state-space coverage (Klissarov et al., 2023), or (3) access to high-level actions (Zhang et al., 2023a; Du et al., 2023). These assumptions are detached from the reality of embodied agents, e.g. in robotics, which don’t come with perfect state or event captioners, pre-existing offline datasets nor with robust, high-level actions. Furthermore, none of these approaches learn an internal model of “interestingness.” Thus, they rely on the LLM, or a distilled module, to continuously guide their exploration.
In this work, we propose SEmaNtically Sensible ExploratIon (Sensei), a framework for Vision Language Model (VLM) guided exploration for model-based RL agents, illustrated in Fig. 1. Sensei starts with a short description of the environment and a dataset of observations (e.g. images) collected through self-supervised exploration. A VLM is prompted to compare the observations pairwise with respect to their interestingness and the resulting ranking is distilled into a reward function. When the agent explores its environment, it receives semantically-grounded exploration rewards from Sensei. It learns to predict this exploration signal through its learned world model, corresponding to an internal model of “interestingness”. The agent improves its exploration strategy by aiming for states for which it predicts a high interestingness and then branching out to uncertain situations. Our main contributions are as follows:
@@ 2 Method
Figure 2: Intrinsic rewards in Sensei: (a) Prior to exploration, we prompt GPT-4 to compare images with respect to the interestingness for a certain environment. From the resulting ranking we distill a reward function $R_{\psi}$ using VLM-Motif. (b) Later, an agent learns an Rssm world model from exploration. From each state the agent predicts two intrinsic rewards, i.e., the distilled semantic rewards $r^{\mathrm{sem}}_{t}$ and uncertainty-based rewards $r^{\mathrm{dis}}_{t}$.
2.1 Reward function distillation: MOTIFate SENSEI
@@ 2.3 Exploration policy: Go and Explore with SENSEI
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: Semantic exploration rewards for example trajectories: From top to bottom we show example trajectories for MiniHack KeyRoom and KeyChest (see Fig. 10 for map view), Robodesk, and Pokémon Red.We showcase rewards from VLM-Motif distilled from GPT-4 annotations. The reward trajectories peak at the “interesting” moments of exploration, such as opening a drawer in Robodesk, picking up the key in MiniHack, or landing a critical hit in Pokémon. For Robodesk we show reward trajectories for both Sensei and a version of Sensei with a more general, zero pre-knowledge prompting strategy (Sensei general, see Sec. 4.2.2).
@@ 3 Related work
Intrinsic rewards are applied either to facilitate exploration in tasks where direct rewards are sparse or in a task-agnostic setting where they help collect diverse data. Many different reward signals have been proposed as exploration rewards (Baldassarre & Mirolli, 2013), such as prediction error (Schmidhuber, 1991; Pathak et al., 2017; Kim et al., 2020), Bayesian surprise (Storck et al., 1995; Blaes et al., 2019; Paolo et al., 2021), learning progress (Schmidhuber, 1991; Colas et al., 2019; Blaes et al., 2019), empowerment (Klyubin et al., 2005; Mohamed & Jimenez Rezende, 2015), metrics for state-space coverage (Bellemare et al., 2016; Tang et al., 2017; Burda et al., 2019) and regularity (Sancaktar et al., 2024). While effective for low-dimensional observations, such objectives are challenging to apply for high-dimensional image observations. Here, alternatives are employing low-dimensional goal spaces (Colas et al., 2019; OpenAI et al., 2021; Nair et al., 2018; Pong et al., 2019; Zadaianchuk et al., 2021; Mendonca et al., 2021) or learning latent world models (Hafner et al., 2019a, 2023; Gumbsch et al., 2024) that can be employed for model-based exploration (Pathak et al., 2019; Sekar et al., 2020). In particular, Plan2Explore (Sekar et al., 2020) uses ensemble disagreement of latent space dynamics predictions as an intrinsic reward. While this is a very general strategy for exploration, this could be limited in more challenging environments where semantically meaningful or goal-directed behavior (Spelke, 1990) is needed for efficient exploration.
@@
Exploration with foundation models: Recent improvements of in-context learning of LLMs open additional ways to explore using human bias of interestingness during exploration (Klissarov et al., 2023; Du et al., 2023; Zhang et al., 2023a) and skill learning (Colas et al., 2020, 2023; Zhang et al., 2023b). Motif (Klissarov et al., 2023) leverages LLMs to derive intrinsic rewards by comparing pairs of event captions, demonstrating its efficacy in the complex game of NetHack (Küttler et al., 2020). Similarly, ELLM (Du et al., 2023) uses LLMs to guide RL agents towards goals that are meaningful, based on the agent’s current state represented by text. Furthermore, OMNI (Zhang et al., 2023a) introduces a novel method to prioritize tasks using LLMs. Thereby, OMNI focuses on tasks that are not only learnable but also generally interesting. LAMP (Adeniji et al., 2023) proposes to use VLMs for reward modulation by first generating a set of potential tasks with an LLM and then generating task-based rewards using VLMs.
@@
Reward-shaping through VLMs: Most works that rely on VLMs as reward sources try to solve the reward specification problem in RL. In these works, a task is assumed to be described as a language caption (Cui et al., 2022; Rocamonde et al., 2023; Baumli et al., 2023; Adeniji et al., 2023), as a goal image (Cui et al., 2022), or as a video demonstration (Sontakke et al., 2023). In particular, RL-VLM-F (Wang et al., 2024) uses a very similar setup to Sensei. Pairs of images from initial rollouts are compared using a VLM to distill a reward function via Motif (Klissarov et al., 2023). However, we assume a model-based setup and do not explicitly prompt the task and distill an environment-specific but general exploration reward.
@@ 4 Results
  4. 4.
    Can Sensei be combined with extrinsic rewards to solve tasks that require substantial exploration?
We answer these questions by (1) illustrating that the semantic rewards obtained from VLM-Motif reflect interesting events in the environment, (2) quantitatively showing that Sensei leads to more interaction-rich behavior during task-free exploration, (3) employing the learned world models to successfully train task-based policies and (4) combining Sensei’s exploration strategy with extrinsic rewards to tackle a challenging environment that cannot be efficiently explored using rewards alone. We use three fundamentally different types of environments:
@@
Pokémon Red is a Game Boy role-playing game where players control a trainer exploring a world of collectible creatures called Pokémon. Exploration is essential for progress, as players must navigate a vast, interconnected world and master a semantic battle system (e.g. Water beats Rock) to defeat Gym leaders and become Pokémon Champion. Our implementation (Whidden, 2023; Suarez, 2024) use the raw game screen as pixel-based observations and discrete actions corresponding to Game Boy button presses (see Suppl. C.3).
In all environments, we collect the initial dataset $\mathcal{D}^{\mathrm{init}}$ with Plan2Explore (Sekar et al., 2020), the current state-of-the art in exploration with pixel-based observations. We collect data from 500k steps in MiniHack and Pokémon Red and 1M steps in Robodesk. For data annotation, we use GPT-4 (details in Suppl. D).
@@ 4.1 Reward function of SENSEI
In Fig. 4 we illustrate how the distilled VLM-Motif reward function $R_{\psi}$ assigns semantic rewards $r^{\mathrm{sem}}_{t}$ for exemplary sequences. In MiniHack, $r^{\mathrm{sem}}_{t}$ clearly jumps for significant events. Frames 2 & 3 in KeyRoom-S15 and KeyChest respectively, are right before the key is picked up. Later, $r^{\mathrm{sem}}_{t}$ increases further once the agent is at the door or chest with a key (Frame 3 in KeyRoom-S15 and Frames 4&5 in KeyChest). For Robodesk, we see that as the robot is interacting with objects, $r^{\mathrm{sem}}_{t}$ increases, e.g., when opening the drawer or pushing the blocks. For Pokémon Red, $r^{\mathrm{sem}}_{t}$ rises while winning a battle, with surges for inflicting damage and drops for setbacks such as getting poisoned. More examples of Robodesk are shown in Suppl. Fig. 17 and examples of Pokémon Red in Suppl. Fig. 26.
@@ 4.2.2 Robodesk
Next, we analyze exploration in the challenging visual control suite of Robodesk. Here we compare 1M steps of exploration in Sensei with Plan2Explore and Random Network Distillation (RND, Burda et al., 2019), a strong model-free exploration approach that uses prediction errors of random image embeddings as intrinsic rewards to maximize state space coverage. Fig. 6 plots the mean number of object interactions during exploration for the three methods. On average, Sensei interacts more with most available objects than the baselines. As a result, in a majority of tasks Sensei receives more task rewards during exploration than Plan2Explore or RND (shown in Suppl. E.3). Qualitatively, we observe that Plan2Explore mostly performs arm stretches [^3] Interestingly, this can still lead to solving tasks. For example, stretching the arm against the sliding cabinet can close it, and stretching the arm toward a block can push it off the table., whereas RND mostly moves the arm around in the center of the screen, mostly hitting buttons, as they are also centered on the table, and occasionally hitting objects.
Thus, our semantic exploration scheme leads to more object interactions than uncertainty-based exploration, even in a low-level motor control robotic environment.
@@
In the previous Sensei experiments, we provided a small environment description in the prompts for the VLM annotations. We investigate whether Sensei relies on this external description in Robodesk, and compare against a version of Sensei using a more general, zero-knowledge prompting strategy (Sensei general). Sensei general first prompts the VLM for an environment description given an image of the environment and uses the generated answer as context to annotate the dataset of preferences (details in Suppl. D.2). As shown in Fig. 6, Sensei general interacts roughly as often with the relevant objects as Sensei, outperforming both Plan2Explore and RND in terms of overall number of object interactions. Thus, injecting external environment knowledge to the prompts is not necessary and this step can be fully automated. This further cements the generality of our approach.
@@
We perform ablations to see (1) how noisy annotations from VLMs affect Sensei compared to an oracle annotator, (2) how much the behavior richness of the initial dataset affects Sensei’s performance (Suppl. Fig. 14), and (3) ablate our Go-explore switching strategy. We observe that as VLMs get better, there is indeed more to gain from Sensei, and richer exploration data helps Sensei bootstrap faster. See Suppl. E.2 for more information. We further showcase the robustness of our Go-Explore switching strategy in terms of hyperparameter sensitivity compared to a variant of Sensei where the semantic and disagreement rewards get fixed weights (Suppl. E.6).
@@ 4.3 Fast downstream task learning
(a) KeyRoom-S15
(b) KeyChest
Figure 7: Downstream task performance in MiniHack: Mean episode scores for (a) KeyRoom-S15 and (b) KeyChest, using world models learned during task-free exploration with Sensei and Plan2Explore (P2X). We also compare learning from scratch with DreamerV3 and PPO. Shaded areas show standard error (10 seeds); curves are smoothed with a window size of 3.
@@
In a supplementary experiment (Suppl. E.7), we analyze fast downstream task learning also on representative Robodesk tasks, and demonstrate more sample-efficient policy learning compared to exploration with Plan2Explore.
  Sensei general  Plan2Explore    DreamerV3
Refer to caption (a) map progress Refer to caption (b) map segments explored Refer to caption (c) highest level of Pokémon in party
Figure 8: Task-based exploration in Pokémon Red comparing Sensei general to Plan2Explore and DreamerV3 for 750k steps. We partition the overall game map into unique map segments for different routes, towns, or buildings (details and full map in Suppl. C.3). We sequentially numbered segments that need to be traversed from game start (0) to the first Gym (9) and plot the percentage of random seeds that reach each segment (a). Temporal exploration trends are visualized by plotting the mean number of unique map segments visited (b) and the highest level of the agent’s Pokémon (c) over episodes, smoothed with a moving average (window size 5). Shaded areas indicate standard error (5 seeds).
4.4 Task-based exploration
Some environments are so complex that meaningful task progress is only possible through effective exploration. We investigate such task-based exploration in Pokémon Red. We compare 750k steps of exploration with Sensei general (using $r_{t}+r^{\mathrm{expl}}_{t}$) to Plan2Explore (using $r_{t}+\beta r^{\mathrm{dis}}_{t}$) and DreamerV3 (using only $r_{t}$). Thus, all agents receive extrinsic task rewards $r_{t}$ during exploration, but only Plan2Explore and Sensei also utilize intrinsic signals.
Progress in the game requires navigating a vast world to reach Pokémon Gyms, as well as assembling and training a strong team to defeat Gym Leaders. To evaluate spatial exploration, we partition the game world into distinct map segments, corresponding to towns, routes, forests, or buildings (details in Suppl. C.3). Over the course of exploration, Sensei consistently discovers new segments, outperforming baselines in terms of total map coverage (Fig. 8(b)). To assess whether this exploration is goal-directed, Figure 8(a) labels the specific segments required to progress from the game start (segment 0) to the first Gym (segment 9), and plots the segments reached per method (high-resolution map in Fig. 12). Only Sensei reaches the first Gym, demonstrating superior exploration aligned with the game’s objectives.
To assess battle-related progress, we track the levels of the agent’s Pokémon (Fig. 8(c)), which serve as a proxy for battle experience and overall strength. Dreamer fails to sufficiently explore the battle system and does not manage to train its Pokémon. Its highest-level Pokémon remains at the same level as at the start of the episode. In contrast, Sensei begins leveling up its Pokémon early during exploration, and from episode 100 onward, it consistently achieves higher levels than Plan2Explore. From episode 390 onward, Sensei, on average, obtains twice as many level-ups per episode as Plan2Explore, indicating greater battle success and a higher potential for future encounters.
Together, these results highlight Sensei’s ability to perform meaningful, goal-directed exploration in rich, open-ended environments. We provide a more detailed analysis in Suppl. E.8.
Second generation of annotations
With more exploration, and due to the vastness of the world in Pokémon Red, Sensei increasingly enters regions outside the distribution of its annotation data ($\mathcal{D}^{\mathrm{init}}$). This leads to degraded semantic rewards. In Suppl. E.9, we show how this limitation can be addressed by refining the reward function through a second round of VLM annotations on Sensei-collected data. When continuing the Sensei run used for annotation for 200 additional episodes, now using the updated reward function, the agent is able to defeat the first Gym and obtain the Boulder Badge. This marks a critical milestone, highlighting the strong potential of iterative semantic reward refinement to unlock meaningful progress in complex environments.